# Conventional and practical metallic superconductivity arising from repulsive Coulomb coupling

### v1

@@ IV Eliashberg-Migdal theory
IV.1 Eliashberg frame-work for Coulomb-induced SC
The role of Coulomb interactions in superconductivity [27] remained only partially understood even after the original BCS theory of superconductivity [17], which was formulated for electrons in momentum space. The central puzzle was how the very weak phonon-mediated attraction can overcome the much stronger Coulomb repulsion [27]. This issue was resolved by considering the retardation in time of the phonon-mediated interaction [28] within the Gorkov-Eliashberg formalism [29]. Intuitively, the idea is that the phonon-mediated attraction is long-ranged in time compared to the Coulomb repulsion, which is instantaneous. This allows electrons to form Cooper pairs by passing through a particular point in space at a later time to avoid the Coulomb interaction. The intriguing consequence of this argument [28] is that a Cooper pairing can occur even from an interaction remaining purely repulsive at all frequencies [28]. However, as apparent from the previous sections and will be discussed in more detail here, the role of Coulomb interactions continues to be poorly understood and calculationally intractable except using drastic and invalid approximations..
The Eliashberg theory mentioned above, which is required to properly account for Coulomb interactions, is characterized by the Eliashberg spectral function [29] that is written as
@@ IV.1 Eliashberg frame-work for Coulomb-induced SC
where $D$ is the Boson propagator and $N(0)$ is the Fermi surface DOS and the integral averages over the Fermi surface. As an aside, it should be noted that the formalism can be generlized to higher angular momentum channels (e.g. KL SC) but including the appropriate form factors in the integral for the spectral function. For the case of pairing by electron-phonon interactions, the above expression can be reformulated [30, 31] in terms of the total screened Coulomb interaction $W$ according to the relation:
@@
where $W(k,\nu)=V_{c}(k)[1-\Pi(k;\nu)V_{c}(k)]^{-1}$ is the RPA screening of the Coulomb interaction, $V_{c}(k)$ is the bare Coulomb interaction and $\Pi(k;\nu)$ is the bare polarizability (which may or may not include phonons depending on the application of interest). The above pairing interaction has been used [32] to predict superconductivity in the electron gas from purely repulsive Coulomb interaction (i.e. without any phonons or other Bosons) treated within the RPA approximation. This was followed by work that augmented the RPA interaction with a Kukkonnen-Overhauser local field corrections [33] to reduce the $T_{c}$ of superconductivity in much of the parameter regime of the electron gas [34, 35, 36] (We note, and as already discussed in Sec. II, we do not believe that using any local field corrections in the theory is an improvement from a fundamental conceptual perspective since such static local field corrections would violate Ward identities and are uncontrolled approximations.). As an aside, we note that the K-L mechanism of superconductivity [1] fits within the RPA framework as well, with the focus shifting to higher angular momentum channels while ignoring the retardation effects. The angular momentum form factors eliminate the repulsive part of the interaction for short-ranged bare repulsion, $V_{c}$, allowing the attractive part to dominate. This section will focus on reexamining the estimated superconductivity in the interacting three-dimensional electron gas.
@@
where $\nu$ is the imaginary frequency and the interaction $W(i\nu)$ has been averaged over the FS in a similar way to the Eliashberg spectral function. Such an averaging over the Fermi surface is justified when the momentum space structure of the Cooper pair is known to be simple. We will elaborate on this further at a later stage. The first term in the above interaction is the Coulomb potential, $W(\infty)=V_{c}$, which is instantaneous and therefore frequency independent. The second term is generated either by some Boson or by screening from other electron-hole pairs. Since Eliashberg theory involves a non-perturbative summation of diagrams, a lynchpin of the theory is provided by Migdal’s theorem [12], which shows that vertex corrections can be suppressed by separating out a low-frequency part from the interaction i.e. $W=W_{<}+W_{>}$ where
@@
would include conventional electron-phonon interactions. Vertex corrections involving $W_{<}(i\nu)$ can be ignored [12] if either one of $\omega_{c}$ or $W_{<}$ is small. Neither is true for the rest of the interaction
@@
which includes the bare Coulomb interaction. The conventional Eliashberg formalism [29] for conventional superconductivity focuses on $W_{<}$ which is used to generate the so-called $\lambda=N(0)W_{<}(0)$, while the high-frequency part of the interaction $W_{>}$ is approximated by a Coulomb pseudopotential $\mu^{*}$ [28], whose value is believed to be $\mu^{*}\sim 0.15-0.2$ for most metals. This leads to a reasonably accurate prediction of $T_{c}$ for conventional superconductors when $\lambda\gtrsim 1$, which are the superconductors with respectable transition temperatures. However, this framework also assumes that the screening of the electron Boson interaction is well-approximated by RPA, which is not necessarily a bad approximation in many cases for nomal metals.
The interpretation of $\omega_{c}$ i.e. the splitting $W=W_{<}+W_{>}$, becomes somewhat more complicated where the pairing arises from electron-electron interactions rather than an external Boson line. The Migdal prescription [12] in the latter case can be viewed as splitting the interaction line in each diagram based on a frequency cutoff, which is not trivial to apply to electron-electron interactions. However, the renormalization group framework [37, 38] maybe viewed as Fermions with a energy $|\omega|<\omega_{c}$ while $W_{>}$ is the Fermion vertex consisting of all Fermion lines with $|\omega|>\omega_{c}$ [38]. This interaction vertex is the building block of the Eliashberg equations [29]. The interaction $W_{<}$ then necessarily contains at least one low-energy electron-hole pair in the Fermi shell. The superconducting transition temperature can be computed from the two-particle irreducible Cooper pair vertex, which be computed in principle irrespective of $\omega_{c}$, though such a calculation would not be subject to perturbation theory. On the other hand, Migdal’s theorem [12] tells us that the Eliashberg equations can be used to determine the two-particle irreducible vertex in the Cooper channel in the energy shell $|\omega|<\omega_{c}$. The choice of $\omega_{c}$ is a balancing act. Choosing $\omega_{c}$ too high introduces uncontrolled errors from ignored vertex corrections. A low choice of $\omega_{c}$ underestimates the transition temperature by restricting the choice of $\Delta$. This is because, as will become clear in the next sub-section the Eliashberg equation for solving $\Delta(\nu)$ can be viewed as a variational minimization where reducing $\omega_{c}$ constrains the set of allowed $\Delta(\nu)$.
@@ IV.2 Solving the Eliashberg equations
Ignoring quasiparticle renormalization (which is likely not critical for this case), the gap equation part of the Eliashberg equation [39] can be written as
@@
where $U$ in the above is the Fermi surface averaged screened Coulomb interaction vertex, which can potentially include vertex corrections [34, 35, 36]. Note that in order to be consistent with Migdal’s theorem [12] as discussed in the last subsection, we will restrict the sum to $|\omega_{m}|<\omega_{c}$. Since superconducting $T_{c}$ can often be exponentially smaller than $\omega_{c}$, the above equation can be difficult to solve. On the other hand, it is possible to find bounds on $T_{c}$ by calculating appropriate integrals following the spirit of the McMillan equation [40]. To do this, we define $\psi_{m}=\phi_{m}/\sqrt{\omega_{m}}$, so that the gap equation can be written in symmetric form
@@
where $\sum_{m>0}\psi_{m}^{2}=1$. This variational form plays a crucial role in this section and is the rationale for being allowed to drop the momentum integrals in the sum. The variational principle allows us to restrict the possible forms of $\phi_{m}$ and estimate a superconducting $T_{c}$, which is a lower bound on the transition temperature. In our case we are restricting $\phi_{m}=0$ to be zero for $|\omega_{m}|>\omega_{c}$ as as well as to be momentum independent over a momentum shell of width $k_{c}$ i.e. $\phi_{m}(k)=\Theta(k_{c}-|k-k_{F}|)\phi_{m}$. This ansatz is included in the fact that $U(\omega_{m})$ is proportional to the screened Coulomb interaction averaged over a shell of width $k_{c}$ around the Fermi surface. Note that the Eliashberg theory with conventional superconductivity [39] usually chooses $k_{c}=0$ because the averaged interaction in the usual electron-phonon case does not have a strong momentum dependence. Unfortunately, this cannot be used for the Coulomb interaction which has divergences at small momenta.
The variational equation Eq. 28 is still quite numerically intensive to solve when the transition temperature is small because it leads to a large density matrix. Therefore, motivated by early works in superconductivity [27, 28, 40] we assume a 2 valued matrix for $\psi_{m}$ i.e. $\phi_{m}=C[A+(B-A)\Theta(\omega_{m}-\omega_{D})]\sqrt{2\pi T}$, where $C$ is a normalization condition determined by the equation
$C^{2}(2\pi T)[A^{2}\sum_{|\omega_{m}|<\omega_{D}}\omega_{m}^{-1}+B^{2}\sum_{|\omega_{m}|>\omega_{D}}\omega_{m}^{-1}]=1.$ (29)
This ansatz was recently compared [41] to more modern renormalization group approaches. This ansatz allows us to write Eq. 28 as
@@
Comparing Eqns. 30 and 31, the coefficients in the above equation are found to be $D_{1}=\int_{\pi T}^{\omega_{D}}\frac{U(\omega-\omega^{\prime})+U(\omega+\omega^{\prime})}{|\omega\omega^{\prime}|}$, $D_{3}=\int_{\pi T}^{\omega_{D}}d\omega^{\prime}\int_{\omega_{D}}^{\omega_{c}}d\omega\frac{U(\omega-\omega^{\prime})+U(\omega+\omega^{\prime})}{|\omega\omega^{\prime}|}$, $D_{2}=\int_{\omega_{D}}\frac{U(\omega-\omega^{\prime})+U(\omega+\omega^{\prime})}{|\omega\omega^{\prime}|}$, $D_{4}=\int_{\pi T}^{\omega_{D}}d\omega/\omega$, and $D_{5}=\int_{\omega_{D}}^{\omega_{C}}d\omega/\omega$, where at finite temperature the integrals represent Matsubara sums according to the convention $\int d\omega\sim(2\pi T)\sum_{m}$. For the purpose of establishing the existence of superconductivity, it will suffice to work with the $T\rightarrow 0$ limit, where the integrals can be taken literally. We will however verify our conclusions against taking the Matsubara sum directly as well.
@@ IV.3 Application to the three dimensional electron gas
We will now apply the fairly general framework above to the three dimensional electron gas within RPA. The starting point for this is the Lindhard polarizability in imaginary frequency [35]
@@
$\left[\frac{1}{2}+\frac{1-z^{2}+u^{2}}{8z}log\frac{(1+z)^{2}+u^{2}}{(1-z^{2})+u^{2}}-\frac{u}{2}tan^{-1}\frac{2u}{u^{2}+z^{2}-1}\right],$ (33)
@@
$\varpi(q,\omega)=-[\frac{1}{2}+\frac{1-z^{2}+u^{2}}{8z}log\frac{(1+z)^{2}+u^{2}}{(1-z^{2})+u^{2}}\\ -\frac{u}{2}tan^{-1}\frac{2u}{u^{2}+z^{2}-1}].$ (36)
@@
$\varpi(q,i\omega)\approx-\frac{1}{2}-\frac{1-z^{2}}{4z}log\frac{1+z}{1-z}+\frac{\pi|u|}{2}.$ (37)
@@
Figure 6: Dimensionless repulsive interaction $U_{0}$ versus the scaled dimensionless pairing interaction $U_{1}\omega_{c}$. The solid line represents the bound for the maximum $U_{0}$ that would support superconductivity. The dots represent the values of $U_{0}$ and $U_{1}\omega_{c}$ (where $\omega_{c}=E_{F}$) for the 3D electron gas for $r_{s}=0.5,0.707,1.0,1.414,2,2.818$ respectively moving from left to right. The momentum shell width for the pair potential is assumed to be $10^{-3}k_{F}$. The electron gas is non-superconducting in this approximation at all the $r_{s}$ values except the last one. The value at $r_{s}=2.8$ superconducts assuming vertex corrections are negligible above $\omega_{c}\sim 0.8E_{F}$ but not for smaller $\omega_{c}$.
@@
$4\{U_{1}(2U_{0}+U_{1}(\omega_{c}-\omega_{D})\}(\omega_{c}-\omega_{D})$
$-U_{0}(\textrm{log}[\omega_{c}/\omega_{D}]+2U_{1}\omega_{D}Log[\omega_{c}/\omega_{D}]))>0.$ (39)
Defining $y=\omega_{c}/\omega_{D}$ provides an upper bound on the repulsive interaction
$U_{0}<\textrm{max}_{y}\frac{U_{1}^{2}\omega_{c}^{2}(1-y)^{2}}{-Log[y]-2U_{1}\omega_{c}(1-y+yLog[y])},$ (40)

### v2

@@ IV Eliashberg-Migdal theory
IV.1 Eliashberg framework for Coulomb-induced SC
The Gorkov-Eliashberg formalism [carbotte1990properties] for calculating $T_{c}$ for electron-phonon superconductivity is one of the success stories of condensed matter physics. However, the role of Coulomb interactions in superconductivity [bogoliubov1960new] remained only partially understood even after the original BCS theory of superconductivity [schrieffer2018theory], which was formulated for electrons in momentum space. The central puzzle was how the very weak phonon-mediated attraction can overcome the much stronger Coulomb repulsion [bogoliubov1960new]. This issue was resolved by considering the retardation in time of the phonon-mediated interaction [morel1962calculation] within the Gorkov-Eliashberg formalism [carbotte1990properties]. Intuitively, the idea is that the phonon-mediated attraction is long-ranged in time compared to the Coulomb repulsion, which is instantaneous. This allows electrons to form Cooper pairs by passing through a particular point in space at a later time to avoid the Coulomb interaction. The intriguing consequence of this argument [morel1962calculation] is that Cooper pairing can occur even from an interaction remaining purely repulsive at all frequencies [morel1962calculation]. However, as apparent from the previous sections and will be discussed in more detail here, the role of Coulomb interactions continues to be poorly understood and calculationally intractable except using drastic and invalid approximations.
The Eliashberg theory mentioned above, which is required to properly account for Coulomb interactions, is characterized by the Eliashberg spectral function [carbotte1990properties] that is written as
@@ IV.1 Eliashberg framework for Coulomb-induced SC
where $D$ is the Boson propagator and $N(0)$ is the Fermi surface DOS and the integral averages over the Fermi surface. As an aside, it should be noted that the formalism can be generlized to higher angular momentum channels (e.g. KL SC) but including the appropriate form factors in the integral for the spectral function. For the case of pairing by electron-phonon interactions, the above expression can be reformulated [kirzhnits1973description, allen1988total] in terms of the total screened Coulomb interaction $W$ according to the relation:
@@
where $W(k,\nu)=V_{c}(k)[1-\Pi(k;\nu)V_{c}(k)]^{-1}$ is the RPA screening of the Coulomb interaction, $V_{c}(k)$ is the bare Coulomb interaction and $\Pi(k;\nu)$ is the bare polarizability (which may or may not include phonons depending on the application of interest). The above pairing interaction has been used [rietschel1983role] to predict superconductivity in the electron gas from purely repulsive Coulomb interaction (i.e. without any phonons or other Bosons) treated within the RPA approximation and in a sense includes pairing mediated by both plasmons and electron-hole pairs. This was followed by work that augmented the RPA interaction with a Kukkonnen-Overhauser local field corrections [kukkonen1979electron] to reduce the $T_{c}$ of superconductivity in much of the parameter regime of the electron gas [buche1990superconductivity, takada1993s, cai2022superconductivity] (We note, and as already discussed in Sec. II, we do not believe that using any local field corrections in the theory is an improvement from a fundamental conceptual perspective since such static local field corrections would violate Ward identities and are uncontrolled approximations.). As an aside, we note that the KL mechanism of superconductivity [Kohn1965] fits within the RPA framework as well, with the focus shifting to higher angular momentum channels while ignoring the retardation effects. The angular momentum form factors eliminate the repulsive part of the interaction for short-ranged bare repulsion, $V_{c}$, allowing the attractive part to dominate. This section will focus on reexamining the estimated superconductivity in the interacting three-dimensional electron gas.
@@
where $\nu$ is the imaginary frequency and the interaction $W(i\nu)$ has been averaged over the FS in a similar way to the Eliashberg spectral function. Such an averaging over the Fermi surface is justified when the momentum space structure of the Cooper pair is known to be simple. We will elaborate on this further at a later stage. The first term in the above interaction is the Coulomb potential, $W(\infty)=V_{c}$, which is instantaneous and therefore frequency independent. The second term is generated either by some Boson or by screening from other electron-hole pairs. Since Eliashberg theory involves a non-perturbative summation of diagrams, a lynchpin of the theory is provided by Migdal’s theorem [Migdal1958], which shows that vertex corrections can be suppressed by separating out a low-frequency part from the interaction i.e. $W=W_{<}+W_{>}$ where
@@
would include conventional electron-phonon interactions. Vertex corrections involving $W_{<}(i\nu)$ can be ignored [Migdal1958] if either one of $\omega_{c}$ or $W_{<}$ is small. Neither is true for the rest of the interaction
@@
which includes the bare Coulomb interaction. The Eliashberg formalism [carbotte1990properties] for conventional superconductivity focuses on $W_{<}$ which is used to generate the so-called $\lambda=N(0)W_{<}(0)$, while the high-frequency part of the interaction $W_{>}$ is approximated by a Coulomb pseudopotential $\mu^{*}$ [morel1962calculation], whose value is believed to be $\mu^{*}\sim 0.15-0.2$ for most metals. This leads to a reasonably accurate prediction of $T_{c}$ for conventional superconductors when $\lambda\gtrsim 1$, which are the superconductors with respectable transition temperatures. However, this framework also assumes that the screening of the electron Boson interaction is well-approximated by RPA, which is not necessarily a bad approximation in many cases for nomal metals. It is clear from the literature of electron-phonon superconductivity that the Eliashberg formalism with RPA screened electron-phonon interactions [carbotte1990properties] with a Coulomb pseudopotential of $\mu^{*}$ and a low Debye frequency clearly is a successful theory of superconductivity. The cutoff $\omega_{c}\sim\omega_{D}$, which is the range of frequency over which the pairing operates in this formalism can be chosen to be much smaller than $E_{F}$ allowing Migdal’s theorem to eliminate vertex corrections [Migdal1958]. In this work, we consider the effect of electron-electron interactions within RPA at larger $\omega_{c}$ to illustrate the role of vertex-corrections play at such $\omega_{c}$.
The interpretation of $\omega_{c}$ i.e. the splitting $W=W_{<}+W_{>}$, becomes somewhat more complicated where the pairing arises from electron-electron interactions rather than an external Boson line. The Migdal prescription [Migdal1958] in the latter case can be viewed as splitting the interaction line in each diagram based on a frequency cutoff, which is not trivial to apply to electron-electron interactions. However, the renormalization group framework [shankar1994renormalization, metzner2012functional] maybe viewed as Fermions with a energy $|\omega|<\omega_{c}$ while $W_{>}$ is the Fermion vertex consisting of all Fermion lines with $|\omega|>\omega_{c}$ [metzner2012functional]. This interaction vertex is the building block of the Eliashberg equations [carbotte1990properties]. The interaction $W_{<}$ then necessarily contains at least one low-energy electron-hole pair in the Fermi shell. The superconducting transition temperature can be computed from the two-particle irreducible Cooper pair vertex, which be computed in principle irrespective of $\omega_{c}$, though such a calculation would not be subject to perturbation theory. On the other hand, Migdal’s theorem [Migdal1958] tells us that the Eliashberg equations can be used to determine the two-particle irreducible vertex in the Cooper channel in the energy shell $|\omega|<\omega_{c}$. The choice of $\omega_{c}$ is a balancing act. Choosing $\omega_{c}$ too high introduces uncontrolled errors from ignored vertex corrections. A low choice of $\omega_{c}$ underestimates the transition temperature by restricting the choice of $\Delta$. This is because, as will become clear in the next sub-section the Eliashberg equation for solving $\Delta(\nu)$ can be viewed as a variational minimization where reducing $\omega_{c}$ constrains the set of allowed $\Delta(\nu)$.
@@ IV.2 Solving the Eliashberg equations
Ignoring quasiparticle renormalization (which is likely not critical for this case), the gap equation part of the Eliashberg equation [marsiglio2008electron] can be written as
@@
where $U$ in the above is the Fermi surface averaged screened Coulomb interaction vertex, which can potentially include vertex corrections [buche1990superconductivity, takada1993s, cai2022superconductivity]. Note that in order to be consistent with Migdal’s theorem [Migdal1958] as discussed in the last subsection, we will restrict the sum to $|\omega_{m}|<\omega_{c}$. Since superconducting $T_{c}$ can often be exponentially smaller than $\omega_{c}$, the above equation can be difficult to solve. On the other hand, it is possible to find bounds on $T_{c}$ by calculating appropriate integrals following the spirit of the McMillan equation [mcmillan1968transition]. To do this, we define $\psi_{m}=\phi_{m}/\sqrt{\omega_{m}}$, so that the gap equation can be written in symmetric form
@@
where $\sum_{m>0}\psi_{m}^{2}=1$. This variational form plays a crucial role in this section and is the rationale for being allowed to drop the momentum integrals in the sum. The variational principle allows us to restrict the possible forms of $\phi_{m}$ and estimate a superconducting $T_{c}$, which is a lower bound on the transition temperature. In our case we are restricting $\phi_{m}=0$ to be zero for $|\omega_{m}|>\omega_{c}$ as as well as to be momentum independent over a momentum shell of width $k_{c}$ i.e. $\phi_{m}(k)=\Theta(k_{c}-|k-k_{F}|)\phi_{m}$. This ansatz is included in the fact that $U(\omega_{m})$ is proportional to the screened Coulomb interaction averaged over a shell of width $k_{c}$ around the Fermi surface. Note that the Eliashberg theory with conventional superconductivity [marsiglio2008electron] usually chooses $k_{c}=0$ because the averaged interaction in the usual electron-phonon case does not have a strong momentum dependence. Unfortunately, this cannot be used for the Coulomb interaction which has divergences at small momenta. This issue is exacerbated by our use of the $T=0$ RPA polarizability function because of which the smoothening of the singularity from thermal broadening is not accounted for in our results. Therefore, our results will only be valid for $k_{c}\gtrsim T/v_{F}$.
The variational equation Eq. 30 is still quite numerically intensive to solve when the transition temperature is small because it leads to a large density matrix. Therefore, motivated by early works in superconductivity [bogoliubov1960new, morel1962calculation, mcmillan1968transition] we assume a 2 valued matrix for $\psi_{m}$ i.e. $\phi_{m}=C[A+(B-A)\Theta(\omega_{m}-\omega_{D})]\sqrt{2\pi T}$, where $C$ is a normalization condition determined by the equation
$C^{2}(2\pi T)[A^{2}\sum_{|\omega_{m}|<\omega_{D}}\omega_{m}^{-1}+B^{2}\sum_{|\omega_{m}|>\omega_{D}}\omega_{m}^{-1}]=1,$ (31)
where $\omega_{D}<\omega_{c}$. This ansatz was recently compared [chubukov2019implicit] to more modern renormalization group approaches. This ansatz allows us to write Eq. 30 as
@@
Comparing Eqns. 32 and 33, the coefficients in the above equation are found to be $D_{1}=\int_{\pi T}^{\omega_{D}}\frac{U(\omega-\omega^{\prime})+U(\omega+\omega^{\prime})}{|\omega\omega^{\prime}|}$, $D_{3}=\int_{\pi T}^{\omega_{D}}d\omega^{\prime}\int_{\omega_{D}}^{\omega_{c}}d\omega\frac{U(\omega-\omega^{\prime})+U(\omega+\omega^{\prime})}{|\omega\omega^{\prime}|}$, $D_{2}=\int_{\omega_{D}}\frac{U(\omega-\omega^{\prime})+U(\omega+\omega^{\prime})}{|\omega\omega^{\prime}|}$, $D_{4}=\int_{\pi T}^{\omega_{D}}d\omega/\omega$, and $D_{5}=\int_{\omega_{D}}^{\omega_{C}}d\omega/\omega$, where at finite temperature the integrals represent Matsubara sums according to the convention $\int d\omega\sim(2\pi T)\sum_{m}$. For the purpose of establishing the existence of superconductivity, it will suffice to work with the $T\rightarrow 0$ limit, where the integrals can be taken literally. We will however verify our conclusions against taking the Matsubara sum directly as well.
@@ IV.3 Application to the three dimensional electron gas
We will now apply the fairly general framework above to the three dimensional electron gas within RPA. The starting point for this is the Lindhard polarizability in imaginary frequency [takada1993s]
@@
$\left[\frac{1}{2}+\frac{1-z^{2}+u^{2}}{8z}\ln\frac{(1+z)^{2}+u^{2}}{(1-z^{2})+u^{2}}-\frac{u}{2}\tan^{-1}\frac{2u}{u^{2}+z^{2}-1}\right],$ (35)
@@
$\varpi(q,\omega)=-\left[\frac{1}{2}+\frac{1-z^{2}+u^{2}}{8z}\ln\frac{(1+z)^{2}+u^{2}}{(1-z^{2})+u^{2}}\right.\\ \left.-\frac{u}{2}\tan^{-1}\frac{2u}{u^{2}+z^{2}-1}\right].$ (38)
@@
$\varpi(q,\omega)\approx-\frac{1}{2}-\frac{1-z^{2}}{4z}\ln\frac{1+z}{1-z}+\frac{\pi|u|}{2}.$ (39)
@@
Figure 9: Dimensionless repulsive interaction $U_{0}$ versus the scaled dimensionless pairing interaction $U_{1}\omega_{c}$. The solid line represents the bound for the maximum $U_{0}$ that would support superconductivity. The circles and squares represent the numerically calculated values of $U_{0}$ and $U_{1}\omega_{c}$ for the 3D and 2D electron gases for $r_{s}=0.5,0.707,1.0,1.414,2,2.818$ respectively moving from left to right. The cutoff $\omega_{c}/E_{F}=1,0.5$ respectively for the 3D and 2D cases. The momentum shell width for the pair potential is assumed to be $k_{c}=10^{-3}k_{F}\sim T_{c}/v_{F}$. The electron gas is non-superconducting in this approximation at all the $r_{s}$ values except the last few. The value at $r_{s}=2.8$ superconducts assuming vertex corrections are negligible above $\omega_{c}\sim 0.8E_{F}$ ($\omega_{c}\sim 0.3E_{F}$ in 2D) but not for smaller $\omega_{c}$.
@@
$U_{1}[2U_{0}+U_{1}(\omega_{c}-\omega_{D})](\omega_{c}-\omega_{D})$
$-U_{0}\ln\frac{\omega_{c}}{\omega_{D}}(1+2U_{1}\omega_{D})>0.$ (41)
Defining $y=\omega_{D}/\omega_{c}$ provides an upper bound on the repulsive interaction
$U_{0}<\textrm{max}_{y<1}\frac{U_{1}^{2}\omega_{c}^{2}(1-y)^{2}}{-\ln{y}-2U_{1}\omega_{c}(1-y+y\ln{y})},$ (42)